OpenStudy (omgthatgirl):

Determine the type and number of solutions of 4x^2-3x+1=0 a. two real solutions <----- my answer b. one real solution c. two imaginary solutions

OpenStudy (anonymous):

what did you get when you factored

rishavraj (rishavraj):

\[D > 0\] - two real and unequal solutions \[D = 0 \] real and equal solutions

rishavraj (rishavraj):

\[D < 0 \] - no real roots

OpenStudy (omgthatgirl):

so its b

rishavraj (rishavraj):

nah....according to ur question D = -7 D is negative

OpenStudy (omgthatgirl):

ok so its a?

rishavraj (rishavraj):

when \[D < 0\] then no real roots..... it means two imaginary solutions/roots

OpenStudy (omgthatgirl):

oh ok thanks

OpenStudy (omgthatgirl):

can you help me with another one

rishavraj (rishavraj):

okay....

OpenStudy (omgthatgirl):

x^2+10x-25=0

rishavraj (rishavraj):

wht u wanna find??

OpenStudy (omgthatgirl):

would this one be two solutions

rishavraj (rishavraj):

find D.....

OpenStudy (omgthatgirl):

d=0

OpenStudy (omgthatgirl):

is that right?

rishavraj (rishavraj):

D = b^2 - 4ac = 100 - 4(-25) (1) = 100 + 100 = 200

OpenStudy (omgthatgirl):

ok pretty confusing

rishavraj (rishavraj):

as \[D > 0 \] it means 2 real and unequal solutions

OpenStudy (omgthatgirl):

so itts two solutions

OpenStudy (omgthatgirl):

@rishavraj

rishavraj (rishavraj):

two real solutions...:)